Hi! Could you help me to solve this one:

m=-12

m/9-30/m
1
Expert's answer
2012-10-08T07:27:40-0400
m=-12

m/9-30/m=|substitute m=-12|=(-12)/9-30/(-12)=-12/9+30/12=-4/3+5/2=(15-8)/6=7/6=1 1/6